# Learning to Partially Defer for Sequences

### v1

@@ 3.1 Surrogate Loss
As described in Section 2.2, minimizing the deferral loss directly is computationally difficult, so identifying an optimal rejector requires proposing and optimizing a convex surrogate loss that is Bayes consistent with respect to ${\cal L}^{\text{Token}}$. To this end, we propose the following surrogate loss function: ${\cal L}^{\phi}(h,r,x,y)=\frac{1}{L}\sum_{j=1}^{L}{\cal L}^{\phi}_{j}(h,r_{j},x,\widehat{y}_{<j},y)$ where
$\begin{split}{\cal L}^{\phi}_{j}(h,r_{j},x,\widehat{y}_{<j},y)=l&\left(y,\widehat{y}_{j}^{h}\right)\phi(r_{j}(x,\widehat{y}_{<j}))\\ &+c_{j}(x,\widehat{y}_{<j},y)\phi(-r_{j}(x,\widehat{y}_{<j}))\end{split}$ (4)
and $\phi$ is a convex binary surrogate loss. $\phi$ could, for example, be the logistic loss, i.e. $\phi(z)=\log\left(1+\exp\{-z\}\right)$ or the square loss, i.e. $\phi(z)=(1-z)^{2}$. ${\cal L}^{\phi}$-risk is characterized by ${\cal R}_{\phi}(h,r)=\mathop{\mathbb{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\phi}(h,r,x,y)]$.
The optimization of ${\cal L}^{\phi}(h,r,x,y)$ is computationally efficient due its convexity. Additionally, the surrogate loss upper bounds ${\cal L}^{\text{Token}}$ upto a scaling factor, providing alignment between reducing the surrogate loss and reducing the desired loss function. The proof of these properties can be found in Section A.1.
To show Bayes consistency of the surrogate, we relate the excess risk using ${\cal L}^{\text{Token}}$ and the excess risk using ${\cal L}^{\phi}$ in Lemma 3.1. The proof employs a specific type of distribution on ${\cal X}\times\{-1,1\}$ with mass on functions of the conditional expectations of $l\left(y,\widehat{y}_{j}^{h}\right)$ and $c_{j}(x,\widehat{y}_{<j},y)$ to relate the two inequalities - a strategy inspired by Mao et al. [15]. The complete proof can be found in Section A.2.
@@
Taking the limit on both sides of the inequality result of Lemma 3.1 proves consistency of the surrogate loss, formally stated in Theorem 3.2. The proof of Theorem 3.2 hinges on the fact that a classification-calibrated surrogate loss satisfies the condition presented in Lemma 3.1. A loss $\phi$ is considered binary classification calibrated if the minimizer of the pointwise $\phi-$risk or $\mathbbm{E}_{y\sim{\cal P}_{{\cal Y}\mid{\cal X}=x}}[\phi(yf(x))]$ always has the same sign as $\mathbb{P}[Y=1\mid X=x]-\frac{1}{2}$. We defer the full proof of Theorem 3.2 and the formal definition of classification calibrated loss to Section A.3.
@@ 3.2 Generalization Bounds
While Bayes consistency is desirable, the property only holds in the limit of infinite data and does not possess finite sample guarantees. In practice, finding a risk minimizer over the entire class of measurable functions is extremely challenging, particularly because ${\cal P}_{X,Y}$ is usually unknown. With access to a rejector training data set ${\cal D}$ with $n$ samples, we can only learn the optimal $\widehat{r}_{n}\in\mathop{\mathrm{argmin}}\limits_{r\in{\cal F}}\widehat{\cal R}_{\phi}(r)$ from a restricted hypothesis class ${\cal F}$ by minimizing the empirical risk with respect to ${\cal L}^{\phi}$$\widehat{\cal R}_{\phi}(r)=\frac{1}{n}\sum_{(x^{i},y^{i})\in{\cal D}}{\cal L}^{\phi}(h,r,x^{i},y^{i})$. We now present an upper bound for the estimation error to characterize the quality of $\widehat{r}_{n}$ with respect to ${\cal F}$. Please refer to Section A.4 for the full proof.
@@
$\begin{split}{\cal R}_{\phi}&\left(\widehat{r}_{n}\right)-{\cal R}^{*}_{\phi}\left({\cal F}\right)\\ &\leq\frac{4\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}\end{split}$
@@ 4 One-time partial deferral
$\begin{split}{\cal L}^{\text{OneTime}}_{\text{Full}}&\left(h,r,x,y\right)\\ &\sum_{j=1}^{L+1}\left(l(y,\widehat{y}_{<j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)=j}\end{split}$ (5)
@@ 4.1 Influence of large $L$
The one-time deferral paradigm is strongly related to the learning to defer with multiple experts framework, which is evident from the similarities between their loss functions (Equations 2 and 5). Each point of the sequence can be seen as corresponding to a distinct “expert” with its prediction consisting of the predictor output up to that point and to the deferred expert’s prediction for the remainder. However, L2D system accuracy tends to plateau with an increasing number of experts, which naturally has ramifications on the cost-effectiveness of the system [35, 36]. In particular, it suggests that allowing for one-time deferral at each point of the sequence may be over-finely resolved, as deferring on neighboring labels may offer a negligible difference in accuracy.
@@
$\begin{split}{\cal L}^{\text{OneTime}}&\left(h,r,x,y\right)\\ &=\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{<j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)=j}\end{split}$ (6)
The ${\cal L}^{\text{OneTime}}$-risk is denoted by ${\cal R}_{\text{OneTime}}(h,r)=\mathop{\mathbbm{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\text{OneTime}}(h,r,x,y)]$.
@@ 4.2 Surrogate loss
Let ${\boldsymbol{g}}$ consist of $|{\cal J}|$ scoring functions $g_{j}:{\cal X}\to\mathbb{R}$ where $j$ indexes the token positions in a sequence and let $r(x)=\mathop{\mathrm{argmax}}\limits_{j\in{\cal J}}g_{j}(x)$. We propose the following surrogate loss function:
${\cal L}^{\psi}$ $\left(h,r,x,y\right)$
$=\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{<j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\boldsymbol{g}}(x),j)$
where $\psi({\boldsymbol{g}}(x),j)$ is a consistent surrogate for the multiclass 0-1 loss or $\mathbbm{1}_{r(x)\neq j}$, and $c_{\text{max}}=\max_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$. Examples of $\psi$ include cross entropy loss, i.e. $\psi_{\text{ce}}({\boldsymbol{g}}(x),j)=-\log\frac{e^{g_{j}(x)}}{\sum_{k\in{\cal J}}e^{g_{k}(x)}}$ or mean absolute error, i.e. $\psi_{\text{mae}}({\boldsymbol{g}}(x),j)=1-\frac{e^{g_{j}(x)}}{\sum_{k\in{\cal J}}e^{g_{k}(x)}}$ [37]. The ${\cal L}^{\psi}$-risk is characterized by ${\cal R}_{\psi}(h,r)=\mathop{\mathbbm{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\psi}(h,r,x,y)]$.
Intuitively, for a given token position $j$, $c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ is large when the subsequence up till $j$ can be predicted correctly and the cost of deferring on the rest of the sequence is small. When this term is large, we desire $\psi({\boldsymbol{g}}(x),j)$ to be small, incentivizing the model to choose $j$ as the deferral point to minimize the surrogate loss. This surrogate loss encourages the rejector to pick a token position which also has minimal costs for deferring on the rest of the sequence - a goal the ${\cal L}^{\text{OneTime}}$ loss function shares.
Just like ${\cal L}^{\phi}$, ${\cal L}^{\psi}$ is also convex and upper bounds ${\cal L}^{\text{OneTime}}$ up to some scale $\gamma$ (See Section B.2 for proof). Theorem 4.1 shows that ${\cal L}^{\psi}$ achieves Bayes consistency.
@@
The proof of the above theorem can be found in Section B.4. If the $l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ terms are equal for all $j$, the weights would be 00. By assuming a non-zero cost difference for at least one pair, we are ensuring that at least one $j$ has a non-zero weight. Such an assumption can be satisfied if there is a sufficient separation in predictive capabilities between the predictor and the expert.
@@ 4.3 Generalization Bounds
Under the empirical risk minimization framework, the optimal rejector $\widehat{r}_{n}\in\mathop{\mathrm{argmin}}\limits_{r\in{\cal F}}\widehat{\cal R}_{\psi}(r)$ minimizes the empirical risk with respect to ${\cal L}^{\psi}$ or ^$\widehat{\cal R}_{\psi}(r)=\frac{1}{n}\sum_{(x^{i},y^{i})\in{\cal D}}{\cal L}^{\psi}(h,r,x^{i},y^{i})$ from a finite hypothesis class ${\cal F}$. Theorem 4.2 presents generalization bounds for ${\cal L}^{\psi}$ (See Section B.5 for the proof).
@@
Suppose $r_{n}\in{\cal F}=\{x\to\mathop{\mathrm{argmin}}\limits_{j\in{\cal J}}g_{j}(x):g_{j}\in{\cal G}_{j}\}$ where ${\cal G}_{j}$ consists of functions having a range of $(-M,M)$ with $M\geq 0$, $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\psi$ is $\rho$-lipschitz continuous with respect to ${\boldsymbol{g}}(x)$ and there exists $u_{\psi}(M)$ such that $\psi\left({\boldsymbol{g}}(x),j\right)\leq u_{\psi}(M)<\infty$ for all $j$ and $r\in{\cal F}$, then with probability $1-\delta$ for a fixed $h$, the following upper bound holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}_{\psi}^{\text{Point}}$:
$\begin{split}{\cal R}_{\psi}&\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\psi}\left({\cal F},h\right)\\ &\leq 2\sqrt{2}q(L)\rho\sum_{j\in{\cal J}}\mathfrak{R}_{n}\left({\cal G}_{j}\right)+u_{\psi}(M)q(L)\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}\end{split}$
@@
Restricting the hypothesis class to have a range $(-M,M)$ practically involves ReLU activation and softmax clipping to prevent the scores ${\boldsymbol{g}}$ from approaching $\infty$- ∞ or $\infty$. For $\psi_{\text{ce}}$, we can expect $u_{\psi_{\text{ce}}}(M)=-\log\frac{e^{-M}}{e^{-M}+Le^{M}}$ and for $\psi_{\text{mae}}$, it is $u_{\psi_{\text{mae}}}(M)=1-\frac{e^{-M}}{e^{-M}+Le^{M}}$. We can further express the upper bound of the excess risk respect to the original token loss ${\cal L}^{\text{OneTime}}$. The characterization of the upper bound and the proof details can be found in Section B.6.
@@ 5 Related Works
Model Cascades [38] share similarities with the learning to defer framework. In cascades, an inference task is progressively handled by a series of models where easier tasks are completed by earlier models in the sequence and complex cases escalate to subsequent models. Similar to confidence-based L2D methods described in Section 2.1, cascading decisions are typically determined by a thresholding deferral rule [39]. Some recent work on applying cascades to large language models [40, 41] has also explored modeling the cascade decision maker. However, model cascades, just like learning to defer methods, ultimately rely on a single model for each prediction. Our work extends cascades by introducing partial deferrals, allowing two models to collaborate on a prediction. Narasimhan et al. [19] also proposes a token-level deferral rule, but their deferral rule compares the predictor’s confidence with expert’s confidence at every token position, which requires many expensive expert queries. Our work learns the expert costs at training time to avoid such expert queries unless the rejector calls for it.
Query Routing methods select a model that provide the highest quality prediction from an ensemble of models. Some query routing algorithms in large language models [42, 43] often require querying either a part or the entire model ensemble to determine the best generation. Some methods train query routing models using specially designed loss functions [44, 45, 46]; these loss functions, however, lack consistency guarantees. Our work not only presents surrogate losses with stronger statistical guarantees but also offers more granular routing decisions.
Many L2D methods draw inspiration from the Cost-Sensitive learning framework. In the example-dependent cost-sensitive setting, each feature $x$ is paired with a cost vector. A cost-sensitive classifier must then learn to identify the index of the cost vector that minimizes the expected cost. Classifiers [47, 48] are often trained with a weighted logistic loss or weighted cross entropy loss function, similar to ${\cal L}^{\text{Token}}$ and ${\cal L}^{\text{OneTime}}$. However, popular methods [49, 50] use losses that are inconsistent with the original cost-sensitive loss.

### v2

@@ 3.1 Surrogate Loss
As described in Section 2.2, minimizing the deferral loss directly is computationally difficult, so identifying an optimal rejector requires optimizing a convex surrogate loss that is Bayes consistent with respect to ${\cal L}^{\text{Token}}$. To this end, we propose the following surrogate loss function: ${\cal L}^{\phi}(h,r,x,y)=\frac{1}{L}\sum_{j=1}^{L}{\cal L}^{\phi}_{j}(h,r_{j},x,\widehat{y}_{<j},y)$ where
${\cal L}^{\phi}_{j}(h,r_{j},x,\widehat{y}_{<j},y)=l\left(y,\widehat{y}_{j}^{h}\right)\phi(r_{j}(x,\widehat{y}_{<j}))+c_{j}(x,\widehat{y}_{<j},y)\phi(-r_{j}(x,\widehat{y}_{<j}))$ (4)
and $\phi$ is a convex binary surrogate loss. $\phi$ could, for example, be the logistic loss, i.e. $\phi(z)=\log\left(1+\exp\{-z\}\right)$, or the square loss, i.e. $\phi(z)=(1-z)^{2}$. ${\cal L}^{\phi}$-risk is characterized by ${\cal R}_{\phi}(h,r)=\mathop{\mathbb{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\phi}(h,r,x,y)]$.
The optimization of ${\cal L}^{\phi}(h,r,x,y)$ is feasible due its convexity. Additionally, it upper bounds ${\cal L}^{\text{Token}}$ upto a scaling factor, providing alignment between reducing the surrogate loss and reducing the desired loss function. The proof of these properties can be found in Section A.1.
To show Bayes consistency of the surrogate, we relate the excess risk using ${\cal L}^{\text{Token}}$ and the excess risk using ${\cal L}^{\phi}$ in Lemma 3.1. The proof employs a specific type of distribution on ${\cal X}\times\{-1,1\}$ with mass on functions of the conditional expectations of $l\left(y,\widehat{y}_{j}^{h}\right)$ and $c_{j}(x,\widehat{y}_{<j},y)$ to relate the two inequalities - a strategy inspired by Mao et al. (2024a). The complete proof can be found in Section A.2.
@@
Taking the limit on both sides of the inequality result of Lemma 3.1 proves consistency of the surrogate loss, formally stated in Theorem 3.2. The proof of Theorem 3.2 hinges on the fact that a classification-calibrated surrogate loss satisfies the condition presented in Lemma 3.1. A loss $\phi$ is considered binary classification calibrated if the minimizer of the pointwise $\phi$-risk or $\mathbb{E}_{y\sim{\cal P}_{{\cal Y}\mid{\cal X}=x}}[\phi(yf(x))]$ always has the same sign as $\mathbb{P}[Y=1\mid X=x]-\frac{1}{2}$. We defer the full proof of Theorem 3.2 and the formal definition of classification calibrated loss to Section A.3.
@@ 3.2 Generalization Bounds
[…]
@@
${\cal R}_{\phi}\left(\widehat{r}_{n}\right)-{\cal R}^{*}_{\phi}\left({\cal F}\right)\leq\frac{4\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$
@@ 4 One-time partial deferral
${\cal L}^{\text{OneTime}}_{\text{Full}}\left(h,r,x,y\right)=\sum_{j=1}^{L+1}\left(l(y,\widehat{y}_{<j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\bm{1}_{r\left(x\right)=j}$ (5)
@@ 4.1 Influence of large $L$
This paradigm closely parallels learning to defer with multiple experts as their objectives are analogous (Equations 2 and 5). Each position $j$ can be viewed as an “expert” whose prediction concatenates the model prefix $\widehat{y}_{<j}$ with the expert-completed suffix $\widehat{y}^{e}_{\geq j}$. Prior work shows diminishing accuracy gains as the number of experts increases, which can reduce cost–effectiveness (Verma et al., 2023; Hemmer et al., 2023). This suggests that allowing deferral at every position may be over-resolved, since adjacent deferral points can yield similar accuracy.
@@
${\cal L}^{\text{OneTime}}\left(h,r,x,y\right)=\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{<j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\bm{1}_{r\left(x\right)=j}$ (6)
The ${\cal L}^{\text{OneTime}}$-risk is denoted by ${\cal R}_{\text{OneTime}}(h,r)=\mathop{\mathbb{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\text{OneTime}}(h,r,x,y)]$.
@@ 4.2 Surrogate loss
Let $\bm{g}$} consist of $|{\cal J}|$ scoring functions $g_{j}:{\cal X}\to\mathbb{R}$ where $j$ indexes the token positions in a sequence and let $r(x)=\operatorname*{arg\,max}_{j\in{\cal J}}g_{j}(x)$. We propose the following surrogate loss function:
${\cal L}^{\psi}\left(h,r,x,y\right)=\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{<j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\bm{g}}(x),j)$
where $\psi({\bm{g}}(x),j)$ is a consistent surrogate for the multiclass 0-1 loss or $\bm{1}_{r(x)\neq j}$, and $c_{\text{max}}=\max_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$. Examples of $\psi$ include cross entropy loss, i.e. $\psi_{\text{ce}}({\bm{g}}(x),j)=-\log\frac{e^{g_{j}(x)}}{\sum_{k\in{\cal J}}e^{g_{k}(x)}}$ or mean absolute error, i.e. $\psi_{\text{mae}}({\bm{g}}(x),j)=1-\frac{e^{g_{j}(x)}}{\sum_{k\in{\cal J}}e^{g_{k}(x)}}$ (Ghosh et al., 2017). The ${\cal L}^{\psi}$-risk is characterized by ${\cal R}_{\psi}(h,r)=\mathop{\mathbb{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\psi}(h,r,x,y)]$.
Intuitively, for a given token position $j$, $c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ is large when the subsequence up till $j$ can be predicted correctly and the cost of deferring on the rest of the sequence is small. When this term is large, we desire $\psi({\bm{g}}(x),j)$ to be small, incentivizing the model to choose $j$ as the deferral point to minimize the surrogate loss. This surrogate loss encourages the rejector to pick a token position which also has minimal costs for deferring on the rest of the sequence - a goal the ${\cal L}^{\text{OneTime}}$ loss function shares.
Just like ${\cal L}^{\phi}$, ${\cal L}^{\psi}$ is also convex and upper bounds ${\cal L}^{\text{OneTime}}$ up to some scale $\gamma$ (See LABEL:sec:proofonetimeconvexupper for proof). Theorem 4.1 shows that ${\cal L}^{\psi}$ achieves Bayes consistency.
@@
The proof of the above theorem can be found in Section B.3. If the $l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ terms are equal for all $j$, the weights would be 0. By assuming a non-zero cost difference for at least one pair, we are ensuring that at least one $j$ has a non-zero weight. Such an assumption can be satisfied if there is a sufficient separation in predictive capabilities between the predictor and the expert.
@@ 4.3 Generalization Bounds
[…]
@@
Suppose $r_{n}\in{\cal F}=\{x\to\operatorname*{arg\,min}_{j\in{\cal J}}g_{j}(x):g_{j}\in{\cal G}_{j}\}$ where ${\cal G}_{j}$ consists of functions having a range of (M,M)(-M,M) with $M\geq 0$, $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\psi$ is $\rho$-lipschitz continuous with respect to ${\bm{g}}(x)$ and there exists $u_{\psi}(M)$ such that $\psi\left({\bm{g}}(x),j\right)\leq u_{\psi}(M)<\infty$ for all $j$ and $r\in{\cal F}$, then with probability $1-\delta$ for a fixed $h$, the following upper bound holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}_{\psi}^{\text{Point}}$:
@@
Restricting the hypothesis class to have a range (M,M)(-M,M) practically involves ReLU activation and softmax clipping to prevent the scores $\bm{g}$} from approaching -$\infty$ or $\infty$. For $\psi_{\text{ce}}$, we can expect $u_{\psi_{\text{ce}}}(M)=-\log\frac{e^{-M}}{e^{-M}+Le^{M}}$ and for $\psi_{\text{mae}}$, it is $u_{\psi_{\text{mae}}}(M)=1-\frac{e^{-M}}{e^{-M}+Le^{M}}$. We can further express the upper bound of the excess risk respect to the original token loss ${\cal L}^{\text{OneTime}}$. The characterization of the upper bound and the proof details can be found in Section B.5.
@@ 5 Related Works
Model Cascades (Viola and Jones, 2001) share similarities with the learning to defer framework. In cascades, an inference task is progressively handled by a series of models where easier tasks are completed by earlier models in the sequence and complex cases escalate to subsequent models. Similar to confidence-based L2D methods described in Section 2.1, cascading decisions are typically determined by a thresholding deferral rule (Yue et al., 2023). Some recent work on applying cascades to large language models (Gupta et al., 2024; Wang et al., 2024) has also explored modeling the cascade decision maker. However, model cascades, just like learning to defer methods, ultimately rely on a single model for each prediction. Our work extends cascades by introducing partial deferrals, allowing two models to collaborate on a prediction. Narasimhan et al. (2024) also proposes a token-level deferral rule, but their deferral rule compares the predictor’s confidence with expert’s confidence at every token position, which requires many expensive expert queries. Our work learns the expert costs at training time to avoid such expert queries unless the rejector calls for it.
Query Routing methods select a model that provide the highest quality prediction from an ensemble of models. Some query routing algorithms in large language models (Jiang et al., 2023; Chen et al., 2023) often require querying either a part or the entire model ensemble to determine the best generation. Some methods train query routing models using specially designed loss functions (Shnitzer et al., 2023; Lu et al., 2023; Ding et al., 2024); these loss functions, however, lack consistency guarantees. Our work not only presents surrogate losses with stronger statistical guarantees but also offers more granular routing decisions.
Many L2D methods draw inspiration from the Cost-Sensitive learning framework. In the example-dependent cost-sensitive setting, each feature $x$ is paired with a cost vector. A cost-sensitive classifier must then learn to identify the index of the cost vector that minimizes the expected cost. Classifiers (Abe et al., 2004; Lin, 2015) are often trained with a weighted logistic loss or weighted cross entropy loss function, similar to ${\cal L}^{\text{Token}}$ and ${\cal L}^{\text{OneTime}}$. However, popular methods (Tu and Lin, 2010; Chung et al., 2015) use losses that are inconsistent with the original cost-sensitive loss.
Selective classification equips a predictor with an abstention option that trades coverage for accuracy. In the noise-free setting, El-Yaniv and Wiener (2010) formalizes the risk–coverage framework. Deep variants implement the selection function via post-hoc confidence thresholding (Geifman and El-Yaniv, 2017) or joint learning(Geifman and El-Yaniv, 2019). Similarly, selective regression extends the reject option to regression (Shah et al., 2022; Jiang et al., 2020; Zaoui et al., 2020). However, unlike learning to defer, these methods treat abstention as a terminal action with a fixed penalty and do not account for the downstream effects of abstention, e.g., accuracy of expert predictions.